Circular No. 7462

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/2000 O1 (KOEHN) B. W. Koehn, Lowell Observatory, has reported his discovery of a

possible comet in the course of the Lowell Observatory Near-Earth-Object Survey. The following CCD observations are available:

2000 UT α_{2000} δ_{2000} m_{1} Observer
July 20.36467 1^{h}03^{m}24.^{s}38 +39^{o}53′01′′.6 17.6 Koehn
20.37584 1 03 24.06 +39 53 03.4  _{”}
20.38693 1 03 23.33 +39 53 04.7  _{”}
20.39803 1 03 23.02 +39 53 07.6  _{”}
22.07410 1 02 18.03 +39 59 08.6  Busch
22.08600 1 02 17.42 +39 59 09.5  _{”}
22.97988 1 01 41.57 +40 02 18.3 17.9 Tichá
22.99126 1 01 41.11 +40 02 20.1  _{”}
23.37434 1 01 25.93 +40 03 38.3  Robinson
23.39039 1 01 24.65 +40 03 41.9  _{”}

B. W. Koehn (Lowell Observatory). 0.59-m LONEOS Schmidt. Apparently a coma 10′′ in diameter and a 50′′ tail in p.a. 80^{o}. Some confusion with surrounding stars.

M. Busch and R. Kresken (Heppenheim). 0.45-m f/4.4 reflector. J. Tichá and M. Tichý (Kleť). 0.57-m f/5.2 reflector. Apparently diffuse;

coma of diameter ∼ 8′′ but no tail. L. Robinson (Olathe). 0.30-m Schmidt-Cassegrain. m_{2} = 18.7.

The following preliminary parabolic orbital elements, by the undersigned, are from nine of the above observations:

T = 2001 Jan. 22.779 TT  ω = 94.^{o}670
Ω = 93.310 }2000.0
q = 5.69793 AU  i = 148.855
2000 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
July 15 1^{h}06.^{m}53 +39^{o}32.^{′}7 5.911 5.862 82.^{o}3 9.^{o}9 17.5
25 1 00.28 +40 09.2 5.732 5.846  91.4 10.0 17.5
Aug. 4 0 52.15 +40 38.1 5.557 5.830 100.6  9.8 17.4
14 0 42.09 +40 56.0 5.392 5.816 109.9  9.4 17.3
24 0 30.21 +40 59.4 5.243 5.802 119.0  8.8 17.2
Sept. 3 0 16.78 +40 44.5 5.117 5.789 127.7  7.9 17.2
13 0 02.28 +40 08.7 5.018 5.776 135.3  7.0 17.1
2000 July 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden